Calculate the molar mass of the following substances.
Sulfur molecule, S8

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

Sulphur molecule, S8 = 8 × 32 = 256 g

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The molar mass of Sulphur molecule S= 8 × Mass of S 
= 8 × 32 = 256 g/mol
 
Hope it helps!
Thank you ✌️